How many gallons of distilled water must be mixed with 5 gallons of 30% alcohol solution to obtain a 25% solution?
Exlpain...

Answer by mananth(16949)

------------------------ percent ---------------- quantity
30% alcohol 30 ---------------- 5 gallons
Water 0 ----------------x gallons
Mixture 25 ----------------5+x gallons

30*5+0x =25(5+x)
150+0 x=125+25x
0x-25x= 125-150
-25x =-25
/ -25
x=1 gallon of Water

m.ananth@hotmail.ca